East Hill Home Healthcare Services was organized five years ago by four friends who each invested $14,000 in the company and, in turn, were issued in total 8,700 shares of $1.00 par value common stock. To date, they are the only stockholders. At the end of last year, the accounting records reflected total assets of $726,000 ($53,000 cash; $511,000 land; $51,000 equipment; and $111,000 buildings), total liabilities of $223,000 (short-term notes payable $120,000 and long-term notes payable $103,000), and stockholders’ equity of $503,000 ($22,000 common stock; $87,000 additional paid-in capital; and $394,000 retained earnings). During the current year, the following summarized events occurred:

  1. Sold 10,200 additional shares of stock to the original organizers for a total of $103,000 cash.
  2. Purchased a building for $65,000, equipment for $13,000, and four acres of land for $21,000; paid $13,000 in cash and signed a note for the balance (due in 15 years). (Hint: Five different accounts are affected.)
  3. Sold one acre of land acquired in (b) for $5,250 cash to another company.
  4. Purchased short-term investments for $19,000 cash.
  5. One stockholder reported to the company that 300 shares of his East Hill stock had been sold and transferred to another stockholder for $3,000 cash.
  6. Lent one of the shareholders $5,100 for moving costs and received a signed, six-month note from the shareholder.


Required:

1. Was East Hill Home Healthcare Services organized as a sole proprietorship, a partnership, or a corporation?

2. During the current year, the records of the company were inadequate. You were asked to prepare the summary of transactions shown above. To develop a quick assessment of their economic effects on East Hill Home Healthcare Services, you have decided to complete the tabulation that follows and to use plus (+) for increases and minus (−) for decreases for each account. The first event is used as an example.

4. Based only on the completed tabulation, provide the following amounts:

5. Compute the current ratio for the current year.Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 4 Required 5 Compute the cuComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 4 Required 5 Based only onRequired 1 Required 2 Required 4 Required 5 During the current year, the records of the company were inadequate. You were ask
Answer #1
Cash Short term
investments
Notes Receivable land Building Equipment Short-term
Notes payable
Long-term
Notes payable
Common Stock Additional paid in capital retained Earnings
Beg 53000 511000 111000 51000 120000 103000 22000 87000 394000
a 103000 10200 92800
b -13000 21000 65000 13000 86000
c 5250 -5250
d -19000 19000
e
f -5100 5100
Total 124150 19000 5100 526750 176000 64000 120000 189000 32200 179800 394000
Total assets = 134800+19000+5700+516500+162000+65000= 915000
Total Liabilities = 120000+189000= 309000
Total stockholders' Equity = 32200+179800+394000= 606000
Cash balance = 124150
Total current assets = 124150+19000+5100= 148250
Current ratio = Current Assets / Current Liabilities
=148250/120000= 1.24